OpenStudy (anonymous):

the volume of a right circular cone is 108pi m^3, and the radius of the base is 3m. Find the central angle of the sector which corresponds to the lateral area of the cone.

OpenStudy (anonymous):

@mathmate @sillynikki1

OpenStudy (anonymous):

@Luigi0210 @Zale101

OpenStudy (badmood):

volume 1/3 * pi * r^2 * h = 108 pi And r = 3 m, so h = 36 m With r and h we can find slanting height L = ./ r^2 + h^2 = 36.13 m Angle of the sector in degree = 2 * r * 180 / L @ = 29.89 degree

OpenStudy (anonymous):

how did you get the h= 36?

OpenStudy (anonymous):

some things error, the answer could be 30.1 degrees, says in the book

OpenStudy (badmood):

yeah my laptop laggs out some times

OpenStudy (anonymous):

angle sector in degrees, is that the working formula?

OpenStudy (anonymous):

how did we get the 30.1 degrees as the answer?